RC Meeting: September, 2020

From clas12-run
Jump to: navigation, search

Back to Run Group F | Daily Run Group F Coordination Meetings


2020

March
19, 20, 21, 22, 23, 24
July
31
August
1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31
September
1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20 21

RC:

Summary of Sep. 3rd

Owl shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3839864
  2. Issues: Beam was finally restored after first tuning on the tagger yoke (harp scans) and then on the FC (more harp scans). Stepan noted an increase in background on BOM and midstream counters again - apparently related to a intermittent vacuum problem around 2H01.
  3. Recorded data: 36.3 M events

Day shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3840091
  2. Issues: Intermittent vacuum problems. RICH recovery didn't work the first time and had been repeated. DAQ crashed (this is happening too often now). Beam off at 3:30.
  3. Recorded data: 30.4 M events

Swing shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3840452
  2. Issues: Beam off until 7:40. Tried to troubleshoot vacuum by temporarily closing valve - did improve vacuum for a time, but then it got worse again.
  3. Recorded data: 19.4 M events

Whole day summary

  1. Total recorded data today: 86 M
  2. Part-II total accumulated data on D2: 1.52 B events - we have exceeded 50% of the expected data for this run!
  3. Concerns: Ongoing vacuum fluctuations (top pressure INCREASING with time). DAQ crashes hard (Cancel-reset-config-download-... does not work) too many times. Upstream Halo counters are high.


Tomorrow daily RC meeting will be at usual time 12:00 PM using the link https://bluejeans.com/7576835804/

Recorded data numbers

Number of events recorded